OpenStudy (anonymous):

Simplify. Write your answers without exponents. (1/32)^-2/5

OpenStudy (anonymous):

\[(1/32)^{-2/5}=\sqrt[5]{32^2}=\sqrt[5]{(2*2*2*2*2)^2}=\sqrt[5]{2^{10}}=\]

OpenStudy (anonymous):

Should be =4

OpenStudy (anonymous):

the fifth root of 32 is 2 because \(2^5=32\) then \(2^2=4\)

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[5]{2^{10}} = 2^{10/5} = 2^2 = 4\] so there are at least 2 ways to do it